OpenStudy (anonymous):

limit of (sin12x/(sin7x)

OpenStudy (kc_kennylau):

(Assume x tends to 0) \[\begin{array}{cl} &\lim_{x\rightarrow0}\frac{\sin12x}{\sin7x}\\ =&\lim_{x\rightarrow0}\frac{12\cos12x}{7\cos7x}\\ =&\frac{12\cos0}{7\cos0}\\ =&\frac{12}7 \end{array}\]